# The Role of Task Complexity in Emergent Abilities of Small Language Models

### v1

@@ 4 Results
Figure 1 shows the transition to learning for all the 15 combinations of the four operations. The bottom plot shows the transition points, found using a logistic fit. The tasks involving MED are colored in green, and tasks involving SUM are colored in red, except for the pure SUM task, which is the curious outlier. We observe a clear hierarchy of tasks: adding MED to the mix makes the task significantly harder, while tasks containing SUM are even harder than MED.
However, surprisingly, the hardest task by far is the pure SUM, requiring at least twice as many parameters as even the mixed SUM tasks. This is peculiar, as one might guess the complexity of the task would depend on the number of tasks mixed together and the hardness of each component task. Again, counter this argument, for mixed SUM tasks we observe the exact opposite: the mixture of all four tasks is easier than mixture of three, which is easier than two. So, something must be happening when we train multiple math tasks jointly, which makes it easier for models to learn. Such behavior has also been reported in prior works on language models lee2024teaching where task diversity is observed to make models smarter. We want to get to the bottom of how this mixture is leading to a dramatic reduction in the learning transition when training on mixed tasks in ListOps.
Figure 2: PCA of embeddings: We choose all models which reached over $90$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure MAX, or all mixtures involving a given operation, e.g. Mixed SUM. Interestingly, pure SUM does not show a discernible structure in the embeddings, whereas all other cases do. Notably, Mixed SUM models exhibit a prominent odd-even separation in PC5.
Embedding Layer.
To understand the difference between the hierarchy of transition points observed, we first examine the embedding layer of the models which learned the tasks (acc. $>90\%$) (see Appendices A.1 and A.2 for more). While the $n_{vocab}\times d_{embed}$ embedding layers of models of different sizes cannot be directly compared, we can compare them using $n_{vocab}$ dimensional principal component analysis (PCA). Moreover, while the PCs of individual models may be noisy, we can find aggregate PCs by first computing the $n_{vocab}\times n_{vocab}$ correlation matrices of a group of models and averaging them. This approach not only reduces the noise but also allows us to combine models of different sizes.
Figure 3: PC5 and 6 of Average Mixed SUM embedding correlations perfectly separate odd and even numbers.
An interesting picture emerges when we average models which learned each individual operation. Figure 2 shows the correlation matrix and top PCs for the embeddings of models trained on the groups we identified in the transition plot: 1) Pure MIN and MAX, 2) Pure and Mixed MED excluding mixture with SUM, 3) Mixed SUM, and 4) Pure SUM. We note that PCs for group 1 (pure MAX or MIN), as well as the mixture groups 2 and 3, exhibit clear, regular patterns. In contrast, the PCs for group 4 (pure SUM) look random.
We also observe that the regular patterns are not all the same. For group 1, 2 and 3, the top three PCs are almost identical, with PC1 always sorting numbers, and PC2 and PC3 being the two lowest frequency waves (1 and 2 peaks). These three are all top eigenvector of the Laplacian of a weighted line graph. Interestingly, PC4 and PC5 differ among these models. In groups 2 and 3, MED and Mixed SUM, PC4 seems close to the next frequency (3 peaks). For PC5, in group 1 it looks random, in 2 it looks like the next wave’s frequency (4 peaks), but for group 3, it does something striking: it identifies parity by separating odd from even numbers. More specifically, a linear combination of PC5 and PC6 does parity (Fig. 3). The presence of the parity mode in the embedding turns out to be a very robust feature of models trained on Mixed SUM tasks. We observe it not only in this grouping, but even very prominently in individual models trained on MAX+MED+SUM (Fig. 4) and mod 10 (Fig. 13, 20). Based on the observations, we pose the following hypotheses:
  1. 1.
    Sorting-based MAX, MIN, MED: PCA in these models show strong features of a line graph where consecutive numbers have edges and have an ordering.
  2. 2.
    Lookup-based Pure SUM: The lack of regularity in PCs suggest a lack of understanding of number properties, hence the algorithm might be lookup and memory-based.
  3. 3.
    Parity and sorting-based solution to Mixed SUM: Parity appears to play a very prominent role in the Mixed SUM algorithm learned.
Considering that much smaller models learned Mixed SUM compared to Pure SUM, we hypothesize that models in these two groups learned entirely different algorithms for SUM. Mixed training seems to lead the model to develop some understanding of number properties, allowing it to use efficient and compact algorithms to solve SUM. Pure SUM models, in contrast, seem to be treating numbers like random symbols. If this is the case, we believe the models need to memorize a sum table, storing $(A,B)\to C$ for $A+B=C$. One way to test the hypothesis of Pure SUM finding symbolic solutions is to design a similar problem where combination of symbols $(A,B)$ turns into $C$, analogous to a shuffled sum table.
Shuffled SUM.
We conduct experiments with a shuffled, symmetric SUM table (i.e. $A+B=B+A$) in Mod 10, one set with Pure SUM and one on MAX+MED+Shuffled SUM. We find that this shuffled version of sum is again more difficult to learn than any of the math operations except for the original Pure SUM, which remains slightly more difficult than even the shuffled version (Fig. 15). Additionally, we observe that MAX+MED+Shuffled SUM is again more difficult than all operations except pure sum, suggesting that the number properties played an important role in the other tasks becoming easier in mixed training. The PCA lacks most of the features we observed in regular numbers, save for partial ordering, as MAX and MED still use the normal number ordering (Fig. 16). Coloring the numbers based on parity, we don’t see the clear parity separation we observe in the regular mod 10 experiments (Fig. 11), as expected.
Figure 4: Evolution of training loss, accuracy, and PCs of the embedding layer, Mod 26: The top shows the evolution of training loss (solid lines) and test accuracy (dashed lines) for models with an embedding dimension of 96 and 3 layers, trained either on pure SUM (blue) or on mixed MAX+MED+SUM (red). Models trained on MAX+MED+SUM were evaluated separately on pure MAX, MED, and SUM subsets; the corresponding accuracy curves are shown in green, purple, and orange, respectively. Curves are the mean on 3 runs, with shaded ±$\pm\sigma$. All models were trained for 20000 iterations. The figures beneath the main plot display PCA embeddings revealing that models trained on MAX+MED+SUM data progressively develop a structured representation of numerical concepts, accompanied by a steady decrease in loss. The also show a prominent parity separation emerging in PC2 and PC4. Parity is colored by red and blue. In contrast, models trained solely on SUM exhibit no clear structure in the embedding space and show long plateaus in the loss curve.
Learning dynamics Pure SUM vs MAX+MED+SUM.
Another piece of evidence for the difference between SUM and the rest comes from the training dynamics. Figure 4 shows the loss curves and the evolution of test accuracy in a pure SUM model vs MAX+MED+SUM, for models with $n_{embed}=96$ and 3 layers, which is slightly above the pure SUM learning transition. The mixed model was evaluated separately for each of the MAX, MED, and SUM test sets to see when each ability emerged. We observed a subtle but important sequence in learning: MAX was learned fastest, followed closely by MED, and then SUM. Crucially, the gaps between the learning of these operations were minimal, with the model beginning to grasp SUM while still perfecting MED. The accuracy curves also show that joint training accelerates the learning of the SUM operation, as evidenced by the much earlier rise of the orange dashed line (joint training) compared to the blue dashed line (pure SUM training), occurring at almost half the number of steps of pure SUM.
Joint Training May Shrink the Search Space.
The model trained on mixed data can take a very different route and converge to a different solution. We see that training on MAX, MIN, or MED all lead to embeddings which exhibit number properties. Because of the CoT steps, the model’s loss decreases when it learns any of the operations involved in the mix. For instance, since learning MAX is much easier than SUM, gradient descent may first learn to solve MAX, which is strongly associated with learning a representation for the digits with the correct ordering. This restriction of the embedding layer can make it significantly easier for the model to learn other number properties and, possibly, find a number-based algorithm for SUM, exploiting the properties of numbers instead of memorizing symbolic patterns.
4.1 Testing the embedding restriction hypothesis.
Figure 5: Learning SUM by finetuning MAX+MED: We train model much smaller than the sum-only learning transition (48 embedding, 2 layers). By switching the training data slowly from MAX+MED to pure SUM (never showing expression mixing all three) the model is able to learn SUM (blue) in this much lower parameter regime. In comparison, the pure SUM models (red) did not learn at this size.
We test the hypothesis that restricting the embedding to what is learned by operations like MAX and MED could lead to smaller models learning SUM. To further explore this hypothesis, we designed a transfer learning experiment: 1) We first trained a model ($n_{embed}=48$ and three layers, which on pure SUM was unable to learn) on MAX and MED operations until proficiency. 2) We then gradually introduced SUM operations, increasing their proportion in the training data from 0% to 100% over 1000 steps, while simultaneously phasing out MAX and MED.
Key findings from this experiment include:
  • The model began learning SUM immediately upon its introduction, despite never seeing mixed expressions (e.g., SUM with MAX or MED).
  • Interestingly, the model started to forget MAX and MED once these operations were no longer present in the training data.
  • Crucially, we verified that the model retained its number-like embedding structure even after MAX and MED were completely phased out.
Perhaps most strikingly, we found that even much smaller models (embedding size 24) could learn SUM perfectly using this hybrid approach, mirroring the efficiency of MAX+MED+SUM models. This resulted in a model capable of performing SUM operations that was 7x smaller than the PURE SUM model, while relying on a more sophisticated, number-like internal representation.
Attention and feedforward layers.
To investigate the internal dynamics of the sum-only vs MAX+MED+SUM (all3) models, we focused on the final layer of a 3-layer transformer network, featuring a single attention head and an embedding dimension of 128 (Appendix E). We find:
  1. 1.
    In the attention sublayer, the MAX+MED+SUM model shows slightly higher ratio values, suggesting that attention mechanisms play a more pronounced role when the model is trained on diverse operations. This could indicate that the attention sublayer is capturing more complex patterns or relationships necessary for handling multiple operations.
  2. 2.
    Conversely, in the feedforward sublayer, the SUM model demonstrates significantly higher ratio values. This suggests that when trained on SUM alone, the model relies more heavily on the feedforward network for computation. This could imply that the SUM operation is being implemented more directly through feedforward transformations.
  3. 3.
    The larger effect size in the feedforward layer (-0.34) compared to the attention layer (0.18) indicates that the difference in behavior is more pronounced in the feedforward component.
These observations suggest a trade-off in how the network allocates its computational resources. The MAX+MED+SUM model appears to leverage its attention mechanism more, potentially to handle the diversity of operations it was trained on. In contrast, the SUM model seems to channel more of its computation through the feedforward network, possibly developing a more specialized but less flexible approach to solving the sum operation.
Figure 6: Attention patterns and layer dynamics in SUM vs MAX+MED+SUM (all3) models. Each panel shows (CoT) solution to a SUM modulo 10 problem, where ’>’ indicates solution steps. The first row shows the input sequence, with curved lines representing attention weights from Layer 2 in a 3-layer network. Black lines highlight attention patterns for a specific digit (shown in orange). Below are shown various layer metrics including the ratio of self-attention to feedforward norms (sa/ffwd), self-attention output norms (sa_out), feedforward output norms (ffwd_out), and ratios of layer outputs to inputs (sa_out/x, ffwd/x). Top: Model trained only on SUM operations shows attention primarily focused on parentheses and structural elements. Bottom: Model trained on MAX+MED+SUM (all3) shows attention strongly connecting to digits being combined in each CoT step, suggesting direct involvement in numerical computation. These distinct patterns suggest fundamentally different algorithms learned by each model.
Figure 7: Comparison of layer output/input ratio distributions for models trained on MAX+MED+SUM vs pure SUM. Left: Attention layer ratio histogram. Right: Feedforward layer ratio histogram. The x-axis represents the log ratio of layer output norm to input norm, while the y-axis shows the density. The black line represents the difference MAX+MED+SUM minus SUM distributions.

### v2

@@ 4 Results
Figure 2 shows learning transitions for single and joint operations. We find that joint training enables models to solve more difficult tasks with fewer parameters and less data. For example, ADD and NADD are the most demanding tasks when trained in isolation, yet when combined with PROD, models that are $\times$2.5 smaller are able to reach 100% accuracy on the task. Similarly, while PROD on prime moduli is challenging on its own, pairing it with MAX reduces the required model size by a factor of $\times$2.6. In contrast, pairing ADD with NADD does not decrease the required model size, nor does pairing PROD-prime with ADD or NADD.
These results challenge the assumption that harder tasks always demand larger models. Joint training helps only when tasks share the right structure: ADD and NADD use similar algorithms but require different representations, so pairing them fails. Meanwhile, PROD with non-prime modulus helps ADD because divisibility patterns create shared computational primitives. The key insight: task synergy requires overlapping low-level building blocks, not superficial similarity. These patterns extend to combinations of 3–4 tasks: joint training on all combinations of MAX, MIN, MED, and ADD (modulo 26) shows the same grouping of transition points by task compatibility (Appendix Fig. 29).
Testing the number properties hypothesis.
We hypothesize that joint training helps because easier tasks reveal number properties (e.g., ordering, parity) that benefit harder tasks. To test this, we designed an experiment that destroys number relationships: we shuffle the addition table so that $A+B=C$ becomes a random mapping with no arithmetic structure (Appendix D). If joint training benefits ADD by revealing number properties, then Shuffled ADD should not benefit from joint training with MAX and MED.
We conducted experiments using a shuffled, symmetric ADD table ($A+B=B+A$) in modulo 26. The shuffled addition proved more difficult to learn than all other mathematical operations except the original ADD (Fig. 3). Crucially, MAX+MED+Shuffled ADD was harder than Shuffled ADD alone: the mixed model reached only 60% accuracy on Shuffled ADD and required twice as many parameters. This is the opposite of what we observe with regular ADD, where joint training helps dramatically.
This confirms our hypothesis: MAX and MED induce number-like embeddings with ordering structure, but Shuffled ADD is incompatible with these representations. PCA analysis shows Shuffled ADD embeddings lack parity separation or ordering (Appendix Fig. 22).
4.1 Embedding Analysis: What Number Properties Emerge?
The Shuffled ADD experiment confirms that number properties are key to joint training benefits. We now examine what specific properties emerge in the learned embeddings. To do this, we analyze the embedding layer using principal component analysis (PCA).
PCA methodology.
In our tokenization scheme, each number and operation is a single token. While embedding layers of different-sized models cannot be directly compared (they have different dimensions), we can compare their $n_{\text{vocab}}\times n_{\text{vocab}}$ cosine similarity matrices. To reduce noise, we average similarity matrices across models that learned each task (accuracy $90\%$), then compute PCA on the averaged matrix. This reveals consistent patterns across model sizes (see Appendices F.4 and F.5 for additional results).
Structured embeddings in joint training.
Figure 4 shows that joint training induces structured embeddings: PROD shows strong parity separation and ordering; ADD+PROD preserves this structure; pure ADD shows weak structure; NADD develops different structure; and ADD+NADD loses structure from both tasks. This explains why ADD+NADD fails despite algorithmic similarity: incompatible representations.
Figure 4: PCA of embeddings reveals structured number representations in joint training. Top 5 principal components of the cosine similarity matrix, colored by parity (yellow=odd, purple=even). PROD: Strong parity separation and ordering structure. ADD+PROD: Preserves PROD’s structure, explaining why joint training helps ADD. Pure ADD: Weak, noisy structure. NADD: Highly structured but different from ADD. ADD+NADD: Loses structure from both tasks, explaining why this pairing fails.
Quantifying embedding structure: separation score.
Beyond parity, we observe that embeddings cluster numbers by their remainders modulo divisors of $n$. To quantify this, we define a separation score: for each candidate divisor $d$, we group numbers by their remainder mod $d$, compute centroids for each group in the top PCs, and measure the ratio of between-group distance to within-group standard deviation. High separation scores indicate strong clustering by that divisor (Fig. 5).
The results reveal: PROD embeddings cluster by divisors of $n$ (e.g., mod-2, 3, 6 for nn=42); prime moduli show no clustering, explaining why PROD-prime is hard; and ADD+PROD inherits PROD’s divisor structure while pure ADD shows weak clustering. This explains why PROD helps ADD only for non-prime moduli: divisibility structure creates transferable primitives.
Figure 5: Embeddings cluster by divisors of the modulus. (a) PCA of PROD-45+ADD-45 embeddings colored by remainder class shows clear mod-3 and mod-5 clustering (divisors of 45). (b) Separation score (between-centroid distance / within-group variance) quantifies clustering strength across moduli. Non-prime moduli (42, 45) show strong clustering by their divisors; prime modulus (43) shows none. ADD+PROD inherits PROD’s divisibility structure; pure ADD shows weak clustering. This explains why PROD helps ADD only for non-prime moduli.
4.2 Learning Dynamics: How Structure Emerges
The embedding analysis shows that joint training produces structured representations. But when does this structure emerge during training? We examine the learning dynamics to understand how easy tasks scaffold hard tasks.
Figure 6: Joint training improves training speed by encouraging structured embeddings during learning dynamics. (Top) Two models are trained: one on a joint MAX+MED+ADD dataset and one on a pure ADD dataset (mod 26). We observe that both training loss (solid) and accuracies (dashed) improve faster for the joint training model across all evaluated tasks (mean $\pm\sigma$ over 3 runs.). In joint training, MAX learns first, then MED, then ADD learns only after the embeddings become structured. (Bottom) PCA snapshots are taken during training (number parity is colored red/blue). Joint training accuracies rise quickly after the embeddings learn structure, which only happens in the joint training model. The pure ADD model shows no embedding structure and exhibits loss plateaus, suggesting it struggles to discover an efficient algorithm.
Figure 6 compares pure ADD vs. MAX+MED+ADD ($n_{\text{embed}}=96$). In joint training, MAX learns first, then MED, then ADD—and structured embeddings emerge before ADD accuracy rises, suggesting structure enables learning. Joint training converges in half the steps; pure ADD shows no structure and loss plateaus.
4.3 Causal Test: Can Learned Structure Transfer?
To establish causation, we test whether structure from easy tasks transfers to hard tasks. We design a curriculum (Fig. 7): train on ADD+PROD, gradually shift to pure ADD, and compare to a baseline trained only on ADD.
The results are striking: baseline ADD stays at $\sim$15% accuracy, but curriculum achieves >>95%. Structured embeddings develop during ADD+PROD and persist through the transition.
Figure 7: Curriculum learning enables transfer of structured representations. Left: PCA of embeddings at 0K, 10K, 20K, and 30K steps (top), colored by parity (red=odd, blue=even). Structure emerges during joint ADD+PROD training (mix ratio rises at 10K) and persists as the curriculum shifts to pure ADD. Loss and accuracy curves (bottom) show the curriculum transition. Right: Baseline ADD stays at $\sim$15% accuracy; curriculum reaches >>95%. Shaded: $\pm 1\sigma$ over runs.
[…]
4.4 Scaling Across Moduli
Do the benefits (and failures) of joint training persist across different moduli? We systematically measure transition points for various task combinations across moduli ranging from 10 to 80 (Fig. 8). The results cleanly separate into cases where joint training helps versus cases where it does not.
Figure 8: Joint training helps only with compatible tasks when scaling across moduli. Top (helps): ADD+PROD (left) reduces transition points by $3.6\pm 1.6\times$ for non-prime moduli, while MAX+PROD (right) reduces PROD transition by $3.3\pm 1.3\times$ even for prime moduli. Bottom (no benefit): ADD+PROD (left) provides no benefit for prime moduli (ratio $0.85\pm 0.25\times$), while ADD+MAX+PROD (right) increases PROD-prime transition (ratio $0.77\pm 0.13\times$). Modulus values are indicated on the x-axes and as numbers on each scattered point.
Joint training helps:
ADD+PROD reduces transitions by $3.6\pm 1.6\times$ for non-prime moduli; MAX+PROD reduces PROD transitions by $3.3\pm 1.3\times$ even for primes (MAX provides ordering structure).
Joint training fails:
ADD+PROD provides no benefit for prime moduli ($0.85\pm 0.25\times$); ADD+MAX+PROD actually hurts PROD-prime ($0.77\pm 0.13\times$). These patterns hold robustly across moduli (Appendix Fig. 44).
4.5 Generalization Beyond Arithmetic: Permutation Groups
To test generality, we extend ListOps to permutation groups with no numerical structure. We construct $6\times 6$ block-diagonal permutation matrices (two $3\times 3$ blocks) with three operations: OP (full matrix, 36 elements), OP_TOP and OP_BOTTOM (individual blocks, 6 elements). The sub-block operations decompose the full operation into simpler components—analogous to how MAX helps ADD.
Figure 9: Joint training benefits generalize to permutation groups. (a) Setup: $6\times 6$ block-diagonal permutation matrices with two $3\times 3$ blocks. OP acts on the full matrix (36 elements); OP_TOP and OP_BOTTOM act on individual blocks (6 elements each). (b) Transition curves: joint training vs. pure OP. (c) Sub-block operations reduce the OP transition point by $5\text{--}6\times$, confirming that decomposing complex operations into simpler primitives helps learning—even in non-numerical domains.
[…]
5 Discussion
Why does joint training help?
[…]
Toward a predictive theory: output distributions.
We hypothesize that task difficulty relates to the output distribution of $\text{Op}(a,b)=c$: tasks with uniform output distributions (like ADD) offer no patterns to exploit, while tasks with peaked distributions (like PROD at non-prime moduli) allow divide-and-conquer strategies—learning frequent outputs first, then rarer ones. This explains why PROD difficulty changes drastically with modulus: at primes, the output distribution is nearly uniform; at non-primes, divisibility creates structure. The parity and modular clustering we observe in embeddings may emerge from such divide-and-conquer: e.g., in ADD20, even+even and odd+odd both yield even results, creating exploitable structure. Our separation score (Fig. 5) provides quantitative evidence: tasks inducing higher separation by divisors tend to help harder tasks more.
Beyond arithmetic.
Our permutation group results suggest these effects extend beyond arithmetic. By Cayley’s theorem, every finite group is isomorphic to a permutation subgroup, so these principles may apply to a large class of algebraic tasks.
Connections to curriculum learning.
Joint training and curriculum learning both exploit task relationships. Our transfer experiments show that sequential curricula benefit from the same mechanism: structured representations learned on easy tasks transfer to hard tasks. This suggests a general principle: pair tasks sharing computational primitives, not merely similar difficulty.
Algorithmic complexity.
Minimum model size depends not just on task complexity but on the algorithmic complexity of the discovered solution—brute-force memorization requires far more resources than efficient computation. Joint training appears to bias models toward efficient algorithms by constraining the representation space.
Broader applicability.
Prior work shows training arithmetic with text improves accuracy (Lee et al., 2024), and structured patterns help downstream tasks (Yin and Yin, 2024; Zhang et al., 2024), suggesting these principles may transfer across domains.
Limitations.
Our synthetic tasks may not directly transfer to natural language with ambiguous outputs, though they suggest principles for curriculum design in pretraining. We focus on small models; effects at larger scales remain untested. Our PCA-based analysis captures only linear structure, and while our output distribution hypothesis provides partial predictions, a complete theory of task compatibility remains future work.
Impact Statement
[…]
Appendix A Related works
Scaling laws for language models characterize how validation loss scales with model size (number of parameters) (Hoffmann et al., 2022), compute (FLOPs) (Muennighoff et al., 2023), dataset size (number of examples) (Hestness et al., 2017), and information capacity (in bits) (Allen-Zhu and Li, 2024), offering a foundation for model design. However, this perspective overlooks critical aspects of learning. Our results show that the composition of the training data—specifically, the mix of tasks—can significantly alter what and how a model learns. Even models with identical sizes can follow different learning trajectories depending on the training setup, suggesting that task structure shapes the internal algorithms that emerge during training.
Joint training Joint training on compositional tasks allows models to acquire the foundational primitives necessary for solving complex operations, often leading to abrupt performance improvements once all constituent skills are learned (Lubana et al., 2024; Okawa et al., 2023). Similarly, training on multiple arithmetic tasks has been shown to improve accuracy on individual operations, highlighting the benefits of shared representations for compositional generalization (Lee et al., 2024).
Number representation Neural networks trained on modular addition tasks develop structured embedding spaces that reflect underlying arithmetic operations. Mechanistic analyses have shown that small models can exhibit ordered number patterns and implement distinct algorithmic strategies, depending on initialization and hyperparameters (Zhong et al., 2023). Some models converge to known solutions such as the Clock algorithm, while others discover novel procedures like the Pizza algorithm, illustrating the algorithmic diversity that can emerge from fixed training data. Periodic structures in the embeddings can be characterized via Fourier analysis, offering additional interpretability (Nanda et al., 2023). These behaviors have also been linked to grokking dynamics, where models abruptly generalize after extended training, accompanied by the emergence of structured embedding patterns (Power et al., 2022; Liu et al., 2022).
In general language models performs poorly on symbolic mathematical tasks (Frieder et al., 2023; Dziri et al., 2023; Dave et al., 2024) such as the ListOps dataset (Nangia and Bowman, 2018) used in this study. Models often struggle with generalization, tending to memorize tasks rather than simulate the underlying algorithms. While mathematical tasks prove challenging for large language models to learn, they provide a controllable playground to test how models learn different tasks and to evaluate their accuracy quantitatively. Furthermore it allows us to tune the task difficulty by combining different operations.
Here, we investigate how language models acquire arithmetic skills by training small models on the ListOps dataset, which enables explicit evaluation through structured mathematical expressions. Adopting a bottom-up approach, we find that models learn to solve these tasks once the number of trainable parameters surpasses a critical threshold. This threshold shifts with the difficulty of the target operation (MAX = MIN << MED << ADD), indicating a dependency on task difficulty. Surprisingly, joint training on multiple operations often facilitates learning, outperforming models trained on individual operations (e.g. MAX+MED+SUM << ADD). This suggests that task diversity can ease optimization by promoting shared representations.